OpenStudy (anonymous):

seven apples and one orange cost 98 cents. two apples and three oranges cost 66 cents. find the cost of one apple and one orange

OpenStudy (anonymous):

i bet you are supposed to solve \[7x+y=98\\ 2x+3y=66\] where \(x\) is the number of apples and \(y\) is the number of oranges

OpenStudy (anonymous):

what next ?

OpenStudy (zzr0ck3r):

I would suggest either 1) multiplying the first equation by \(-3\) and then adding the two equations. 2) Solve the first equation for \(y\) and plug that into the second equation.

OpenStudy (radar):

Actually x is the cost of one apple and y is the cost of 1 orange.

OpenStudy (radar):

I would use substitution to solve. Take the first equation 7x + y = 98 and express y in terms of x getting: y = -7x + 98 Did you follow that with understanding?

OpenStudy (radar):

It appears that you went off line so I would assume you can take it from here.

OpenStudy (radar):

It is now simply a matter to substitute for y in the second equation and solve for x, then nail y.